It's still 1 in 6. It's always 1 in 6. Each roll is independent.
The chance is 1/6 as always.
What is the probability of flipping a coin 10 times and landing on heads each time?
If it is a fir coin, the probability is (1/2)10 = 1/1024.
How is the Pythagorean theorem used to develop the equation of a circle in standard form?
The Pythagorean theorem is used to develop the equation of the circle. This is because a triangle can be drawn with the radius and any other adjacent line in the circle.
What is the probability of dying by shark attack?
First of all, I want you to excuse my English that is not as good as i would like.
Youhave more chances to win at the lotery
You have more chances to die:
hitten by lighnting (200 deads each year)
crossing the street
hit by a coconut falling from a palm(more than 100 deads each year)
Each year, the Bengal tiger kills 200 persons in Pakistan and India
There are between 20-30 shark attacks each year and the half of this survive so...be affraid of several things around you but not of sharks.
Hugues from Québec
Actually the 20-30 attacks is a false statistic I mean I go swiming to New Smyrna beach and it's the shark attack capital of the world. the number of attacks their for 2008 is 19 so I doubt the rest of the world is only 11 but it is tru that all these other deaths are more likely
The odds of dying to a shark attack are about 1 in 3,700,000. Both of the posters above are correct in that there are more important things to worry about. The odds of dying due to accidental drowning and submersion during your lifetime are 1 in 1,123 according to the National Safety Council. While sharks are more scary, riptides are what you should really be aware of.
What is the probability of rolling a 7 with two dice?
The probability or rolling a 7 with one roll of a pair (2) dice is 0.166666.... We get this number by dividing 6 by 36 (or 1 by 6). Let's review what a probability is and then find out how we get the 6 and the 36.
We know that probabilities range from zero (0) to one (1), inclusive. A probability of 0 means that a given result cannot happen. Like rolling a 13 with one toss of a pair of dice. The largest number that can appear is 12, so rolling a 13 cannot be done. The probability of that happening is 0.
A probability of 1 means it must happen. Like rolling a number between 2 and 12, inclusive, with one roll of the dice. Rolling a 2, 3, 4, 5, 6 7, 8, 9, 10, 11, or 12, is something that must happen, so the probability of that is 1. All probabilities that something will happen will fall on or between 0 and 1. All of them. Some folks often interchange the term probability with "the odds" of something happening, and this might be the cause of some confusion. Now that we've reviewed probability, let's look at odds and see where the numbers come from.
There are 6 ways to make a 7 with two dice. They are 1-6, 2-5, 3-4, 4-3, 5-2, and 6-1. And there are 36 (6 x 6) possible outcomes with a single roll of two dice. The odds of a 7 coming up are 6 in 36, or 1 in 6. As this is true, the probability of the 7 appearing will be the number of times the combination is possible (6) divided by the total possible combinations (36), which leads us to 6 divided by 36 or 1 divided by 6, or 0.166666....
Use the link to the Related question to see about the odds of rolling that 7 with two dice.
Write an expression equal to 100 that uses seven 4's and the addition sign?
44 + 44 + 4 + 4 + 4 = 100
Sorry to break everyone's heart and soul but not everything is possible. You can't slam a revolving door while it's in motion. You can't grow wings and fly.
So, I have one last thing to say to all of you:
"Nothing is impossible, the impossible just takes infinitely longer."
What are the advantages of stratified random sampling?
There are many advantages of using the stratified random sampling. Some of them are, ability to reduce human potential in choosing the cases in sample, statistical conclusion fro data collected, improving representation of strata etc.
with love,
Good morning or whatever the weather may be,my name is alisa I'm tall and nice looking girl i just decided to drop you some words just to say and how was today, i will like to known more about you, and also i will like to tell you about me,please i will be very happy if you can reply me so that we can go further to known each other.
thanks I'm waiting for your reply,
yours Alisa.
(alisa4u2009@yahoo.ca)
0.97
Can you use a tree diagram to calculate the probability that the sum of the dice is an even number?
Yes, it is possible, but it is far easier to simply sit and think about the problem for a bit. For the sum of two numbers to be even, their parities must be the same (i.e. they must either be both even or both odd).
If you roll a die, and it comes out to be even, then the other die must also be even for the sum to be an even number. The same is true if the die comes out to be odd. Then you can make sort of a generalized tree diagram, where you group outcomes together. Then the two possible outcomes for a die are even and odd. If you construct a tree diagram, it will split into two branches, with one even and the other odd. The same will occur for each of these branches, with a total of four branches at the end. Half of them will satisfy your requirement; that is, in half of all outcomes the sum of the dice is an even number.
Sorry for the rather unnecessarily verbose answer.
What is the least number that has three different prime factors?
If by "least number" you mean "smallest positive integer", then the answer is the product of the three smallest primes: 2x3x5 = 30
The chance it is red is 1/2 The chance it is a jack is 4/52 Now add the two together: 1/2 +4/52 = 30/52 However this double counts the jacks that are red 2/52 so 30/52 - 2/52 = 28/52 Therefore the probability you draw a jack or a red is 7/13
If you flip a coin twice what are the chances that you will get one head and one tail?
One in four or one in two, depending on whether or not the order for heads-tails is important
If you flip a coin once what is the set of possible outcomes for this?
enless you include it landing on it's side the two possible outcomes for this are: Heads and Tails
A number cube is rolled three times. What is the probability of rolling a 2 each time?
the chances of rolling a 2 each time during 1 set of 3 rolls is -1 in 216- sets. i believe the formula is: 1 in 6 - the chance of rolling a"2" then multiply the chances for each separate roll 1/6 x 1/6 x 1/6 = 1/216
If there are 10 people in a room what is the probability that any 2 people have the same birthday?
11.7%. (But when there are 23 people, the odds are 50.7%.) In order for 2 people to have the same birthday, then there can only be 9 different dates out of 365 (actually 37 out of 1461). This is roughly equivalent to 2.53% of possible dates. However, the chance of having matching birthdays to any other birthday is actually much larger, because EACH individual has up to 9 chances out of 365 to match another. The probability is calculated using the product of the series p(n) = 1-(n-1/365), which compares the likelihood of all possible pairs of identical birthdays. The chance that no two share a birthday (including the leap day) can be calculated as the product of the terms n = 1 to 10 of [(1461-(4n+1)/1461]. This should also equate to a likelihood of 88% because the Leap Day is only 1 chance in 1461. (see related link)
What is the probability of Down syndrome if the parents are Asian?
the probobility for downsyndrome depends if any of the parents had it or if a relative had it. it doesn't matter if you are Asian or not. you might get it or you may not. it all depends if you or your relatives have it.
What is the probability that exactly 3 of the 10 are left handed?
very high, that is only slightly higher than average.